OpenStudy (anonymous):

I am trying to solve this system of equations (i will type it in a bit) I can't seem to find the answer I need to complete this problem

OpenStudy (anonymous):

alpha, theta, L1,L2, L3 and y are known. \[L_{1}\sigma \cos \theta-L_{2}\phi \sin \alpha=0\] \[x-L_{1}\sigma \sin \theta +L_{3}\phi \cos \alpha=0\] \[L_{1}\sigma \cos \theta + L_{3}\phi \cos \alpha -y=0\] There are 3 equations and 3 unknowns.. this should be "do-able" I need ti find x in terms of y basically..